OpenStudy (anonymous):

Determine the sum for the arithmetic series: n = 41, t1 = 17, t41 = 245 A. 303 B. 5,371 C. 10,824 D. 7,105

OpenStudy (anonymous):

A(n) = A(1) + a (n-1) A(41) = A(1) + a (41-1) a = (245 -17)/40 a = 5.7 A(n) = 17 + 5.7(n-1) or = 11.3 + 5.7n S(n) = ∑ (11.3 + 5.7i) = 11.3 n + 5.7 (½) (n)(n+1) S(41) = 11.3 (41) + 5.7 (½) (41)(42) = 5371

OpenStudy (anonymous):

Thanks so much!

OpenStudy (anonymous):

your welcome again =)

OpenStudy (anonymous):

Lol :)

OpenStudy (anonymous):

I have another question for you